What is the range of y=sqrt(16-x^2)
Answer by stanbon(57387)

What is the range of y=sqrt(16-x^2)
----
Sqrt can never be negative.
Solve 16-x^2 >= 0
(x-4)(x+4) <= 0
Draw a number line and plot x = -4 and x = 4
----
Test a value in each of the three intervals:
(x-4)(x+4) < 0
---
If x = -10 ; -*- < 0 ; false
If x = 0 ; -*+ < 0 ; true, so solutions in [-4 , 4]
If x = +10 ; +*+ < 0 ; false
=====================================
Range: -4 <= x <= 4
========================
Cheers,
Stan H.